There are two complementary angles. The larger angle is 15 degrees more than twice the smaller angle. Fine the measure of each angle.
Answer by nerdybill(7384)

There are two complementary angles. The larger angle is 15 degrees more than twice the smaller angle. Fine the measure of each angle.
.
The key is to remember that two angles that are complementary add up to 90 degrees.
.
Let x = smaller angle
then
2x+15 = larger angle
.
x + 2x+15 = 90
3x+15 = 90
3x = 75
x = 25 degrees (smaller angle)
.
larger angle:
2x+15 = 2(25)+15 = 50+15 = 65 degrees